if f, g, and h are functions, and m and b are real numbers such that h(x)=(1-x)g(x)+2=(2-x)f(x)-1=mx+b, for all x, find the values of m and b.
Answer by ikleyn(53751)

.
if f, g, and h are functions, and m and b are real numbers such that h(x)=(1-x)g(x)+2=(2-x)f(x)-1=mx+b, for all x, find the values of m and b.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


Use punctuation marks accordingly and properly.